OpenStudy (anonymous):

the sums of a two digit number is 6. when the digits are reversed the new number is 36 more than the original number. find the original number

OpenStudy (nurali):

Let x and y are two digits of a number number = n Sum of the digits x+y=6.....................(1) When x is ten and y is unit 10x+y=n.............(2) When the digits are reversed, the new number is 36 more than the original number Now y is ten and x is unit 10y+x=n+36.............(3) Subtract (2) from (3) 10y+x=n+36.............(3) -10x-y=-n.............(2) -9x+9y=36 9(y-x)=36 9(y-x)/9=36/9 y-x=4...................(4) Add (1) and (4) y-x=4...................(4) x+y=6.....................(1) 2y=10 2y/2=10/2 y=5 Put the value of y in (1) x+5=6.....................(1) x=6-5 x=1 Number = 15 When the digits are reversed = 51 (15+36=51)